### v1

# Volatility Spillovers and Interconnectedness in OPEC Oil Markets: A Network-Based log-ARCH Approach

Volatility Spillovers and Interconnectedness in OPEC Oil Markets: A Network-Based log-ARCH Approach
Fayçal Djebari Department of Economics, University of Bejaia, Algeria. Email: faycal.djebari@univ-bejaia.dz;
ORCID: 0009-0002-9265-9541
   Kahina Mehidi Department of Economics, University of Bejaia, Algeria. Email: kahina.mehidi@univ-bejaia.dz    Khelifa Mazouz Department of Economics, University of Bejaia, Algeria. Email: khelifa.mazouz@univ-bejaia.dz;
Cardiff Business School, Cardiff University, United Kingdom. Email: MazouzK@cardiff.ac.uk
   Philipp Otto School of Mathematics and Statistics, University of Glasgow, United Kingdom. Email: Philipp.Otto@glasgow.ac.uk; ORCID: $0$-0002-9796-6682
(July 20, 2025)
This paper examines several network-based volatility models for oil prices, capturing spillovers among OPEC oil-exporting countries by embedding novel network structures into ARCH-type models. We apply a network-based log-ARCH framework that incorporates weight matrices derived from time-series clustering and model-implied distances into the conditional variance equation. These weight matrices are constructed from return data and standard multivariate GARCH model outputs (CCC, DCC, and GO-GARCH), enabling a comparative analysis of volatility transmission across specifications. Through a rolling-window forecast evaluation, the network-based models demonstrate competitive forecasting performance relative to traditional specifications and uncover intricate spillover effects. These results provide a deeper understanding of the interconnectedness within the OPEC network, with important implications for financial risk assessment, market integration, and coordinated policy among oil-producing economies.
Keywords: Oil price volatility, OPEC, Volatility spillovers, Network models, Spatiotemporal econometrics
JEL Classifications: C33, C58, D85, Q43, Q47
1 Introduction
The Organization of the Petroleum Exporting Countries (OPEC) plays a central role in stabilizing global oil markets by coordinating production and pricing among its members. Acting as a cartel, OPEC restricts oil supply to support prices (Kilian, 2009; Al Rousan et al., 2018). Despite this coordination, significant differences persist across member countries regarding production behavior and pricing dynamics. Understanding how volatility transmits across these economies remains an important, unresolved question.
Recent studies have increasingly adopted network-based methodologies to explore volatility spillover and interdependencies. These include dynamic production networks (Al Rousan et al., 2018), time-varying spillover indices (Ji et al., 2018), multilayer spillover structures (Dai et al., 2023), and global energy networks (Uddin et al., 2023). Many of these rely on the connectedness framework of Diebold and Yilmaz (2009, 2012, 2014), which measures directional spillovers using forecast error variance decompositions. However, these approaches often depend on external explanatory variables or impose structure through rolling-window VARs, and do not directly model the volatility process or fully account for the network-based structure of conditional heteroskedasticity.
This paper addresses this gap by asking: Can volatility spillovers among OPEC countries be better captured by embedding network structures directly into volatility models? Our motivation stems from the need for models that reflect both the time-varying nature of volatility and the structural interdependence among oil-exporting countries. In particular, the OPEC framework presents a unique case in which policy coordination coexists with heterogeneity in market responses. This calls for a modeling strategy capable of uncovering latent transmission channels that are both temporal and cross-sectional.
To address this, we develop a network-based volatility modeling framework. Building on the spatiotemporal ARCH modeling literature, such as Otto et al. (2024a), we introduce a methodology that embeds time-series clustering-derived network structures into the volatility specification. This allows us to model volatility transmission not only over time but also across a network of interdependent economies. Our approach does not rely on external variables; instead, it infers spillover dynamics directly from the historical log return data.
We focus on six OPEC countries, selected based on the completeness of their monthly historical log return data from the official OPEC database. We construct several dissimilarity-based weight matrices using raw data Euclidean distances, correlations, and model-implied metrics from CCC-GARCH (Bollerslev, 1990), DCC-GARCH (Engle, 2002), GO-GARCH (Van der Weide, 2002), and the financial networks considered in the network log-ARCH models of Mattera and Otto (2024). These weight matrices define the network structures used in the log-ARCH specification. The model is implemented within a rolling-window forecasting scheme to evaluate out-of-sample performance.
Our empirical results show that network-based log-ARCH models incorporating CCC, DCC, and GO-GARCH-derived weight matrices demonstrate strong forecasting performance relative to standard alternatives and substantially reduce the computational complexity associated with the standard multivariate GARCH models. These models yield lower Root Mean Squared Forecast Errors (RMSFE) and Mean Absolute Forecast Errors (MAFE) than most competing specifications. More importantly, they uncover richer and more detailed spillover patterns across the OPEC network. These patterns often align with known geopolitical and institutional ties, even though no such data are explicitly included in the model.
In addition to improving forecast accuracy, the proposed framework enhances interpretability. This reveals the latent transmission channels of volatility that reflect the evolving relationships between member states. For investors, this provides a tool for improved volatility forecasting and dynamic risk management. For policymakers, the model offers early warning signals of breakdowns in coordination or emerging systemic risks. Since the method relies only on return data, it enables real-time assessment without the need for external indicators.
The remainder of this paper is organized as follows. Section 2 reviews the relevant literature on oil price volatility, forecasting, and network-based modeling. Section 3 describes the dataset and presents the key descriptive statistics. Section 4 provides an overview of univariate and multivariate volatility models. Section 5 outlines the methodological framework for the network-based log-ARCH models and the construction of the network weight matrices. Section 6 presents the empirical findings, including the estimation results, network structures highlighting interconnectedness and spillover effects, and the comparative forecasting performance of the proposed models.
2 Literature Review
Since its establishment in Baghdad in 1960, OPEC has aimed to coordinate and harmonize the petroleum policies of its member countries to ensure stable oil prices and adequate returns on investment (OPEC Statute, ). Over the decades, the organization has been shaped by a series of major historical and geopolitical events (OPEC Brief History, ), many of which have had far-reaching effects on the global economy (Jo, 2014). These events have exposed the persistent susceptibility of oil-exporting countries to external shocks, highlighting the complexity of managing coordinated policies in the face of heterogeneous economic responses (Liu et al., 2019). Geopolitical risks, particularly in oil-producing regions such as the Middle East and North Africa, have consistently driven oil price volatility (Mignon and Saadaoui, 2024). For instance, regional tensions and conflicts, along with global crises—such as the Chinese economic slowdown (Hamilton, 2014)—have led to sharp fluctuations in oil prices. More recently, the COVID-19 pandemic amplified this volatility by triggering both supply- and demand-side shocks (Bourghelle et al., 2021; Le et al., 2023). Such compounded disruptions underscore the importance of understanding how volatility transmits across oil-exporting economies, particularly within policy-coordinated networks such as OPEC, where shocks may propagate unevenly across members.
Economies that are heavily dependent on oil exports face considerable challenges due to persistent shocks and uncertainties in global oil markets. These fluctuations hinder long-term economic planning, disrupt investment strategies, and constrain economic growth (Sachs and Warner, 1995). To mitigate these adverse effects, it is crucial for such countries to diversify their energy sources and reduce their reliance on oil revenues, thereby limiting their exposure to oil price volatility (Kilian, 2008, 2014). Diversification enhances economic resilience by insulating national economies from the destabilizing effects of external shocks.
Oil price volatility remains a central concern in economic and financial forecasting, given its influence on key macroeconomic indicators, such as investment, durable goods consumption, and aggregate output. Uncertainty generated by oil price fluctuations can lead to reduced economic activity, particularly in resource-rich exporting nations (Finn, 2000; Kilian, 2008; Elder and Serletis, 2010; Vo and Le, 2021). These countries, as emphasized by Cappelli et al. (2023), are disproportionately affected due to their strong fiscal reliance on oil revenues. Despite decades of research, accurately forecasting oil price movements remains a formidable task due to complex interactions among market agents (Baumeister and Kilian, 2016a).
Volatility in oil markets also has profound implications for corporate investment behavior. Elevated uncertainty about future energy input costs, as discussed by Pindyck (1991), deters firms from undertaking irreversible investment projects. This effect is more pronounced in oil-exporting economies, where future profitability is directly tied to oil price movements. Numerous empirical studies corroborate the negative relationship between oil price uncertainty and investment levels (Henriques and Sadorsky, 2011; Ratti et al., 2011; Wang et al., 2017). The exceptional volatility of oil and gas prices, surpassing that of most commodities, is attributed to supply-demand dynamics, macroeconomic fluctuations, and geopolitical tensions, as evidenced by research papers such as (Regnier, 2007) and (Demirbas et al., 2017).
Conventional econometric models often fail to fully capture the complexity of oil price volatility, particularly its dynamic and interconnected nature within coordinated economic blocs. In response, Engle (1982) and Bollerslev (1986) developed the ARCH and GARCH models, which model volatility clustering in financial time series. These nonlinear frameworks remain foundational in financial econometrics, especially through multivariate extensions such as CCC-GARCH (Bollerslev, 1990), DCC-GARCH (Engle, 2002), and GO-GARCH (Van der Weide, 2002), which accommodate dynamic correlations in the return series. More recently, attention has shifted to models that incorporate spatial and network-based dependencies to account for cross-country volatility transmission. Spatiotemporal ARCH (spARCH) models incorporate contemporaneous spatial and temporal dependence to better capture volatility interlinkages (Otto et al., 2018), while spatial and spatiotemporal log-ARCH models (Sato and Matsuda, 2021; Otto et al., 2024a) and network log-ARCH models (Mattera and Otto, 2024) extend this framework to account for multiplicative effects. A comprehensive synthesis of these advanced approaches is presented in Otto et al. (2024b).
Empirical studies have shown strong interdependencies between oil prices and financial markets (Awartani and Maghyereh, 2013; Guhathakurta et al., 2020; Guan et al., 2024). For instance, Nguyen et al. (2020) document the short-run global repercussions of oil price shocks, while Creti et al. (2014) observe heightened oil–stock market co-movements in oil-exporting countries. However, existing studies often rely on models that assume static or exogenous relationships, limiting their ability to reflect the endogenous and evolving nature of the interdependencies among oil-exporting countries. In response, recent advances—such as time-series clustering methods (Maharaj et al., 2019) embedded into Dynamic Spatiotemporal ARCH models (Otto et al., 2024a)—offer promising tools for inferring network structures directly from return data. This opens new avenues for modeling volatility transmission in a way that reflects both temporal dynamics and cross-sectional spillovers, which is particularly relevant for OPEC economies, where coordinated production masks heterogeneous national responses. By integrating clustering-derived networks with advanced volatility models, our approach captures the latent and evolving interdependencies that shape volatility spillovers in oil-exporting nations. These methods move beyond traditional frameworks by eliminating the need for external variables, allowing for real-time analysis rooted in the structure of the return data itself. This offers both improved forecast accuracy and deeper insights into the underlying transmission channels relevant for policy coordination and risk management.
3 Data, Descriptive Statistics and Preliminary Analyses
The empirical analysis in this study is based on monthly crude oil price data for six key OPEC member countries: Algeria (Saharan Blend), Iran (Heavy), Libya (Es Sider), Nigeria (Bonny Light), Saudi Arabia (Arab Light), and the United Arab Emirates (Murban). These countries were selected based on the availability of a complete and consistent historical time series, which offered the same number of monthly observations across the entire sample. Crude oil prices were retrieved from the official OPEC database and subsequently converted into monthly logarithmic returns. The dataset spans the period from January 1983 to December 2022.
Figure 1 presents the time series of the logarithmic returns and their squared values, capturing the volatility dynamics of the OPEC crude oil markets. Notable spikes in volatility coincide with major global events and financial disruptions, including the 1986 oil price collapse, the Gulf War in 1990, the 2008 global financial crisis, the 2014 oil price downturn (Baumeister and Kilian, 2016b; Prest, 2018), the COVID-19 pandemic in 2020 (Bourghelle et al., 2021; Le et al., 2023), and the Russia–Ukraine conflict (Zhang et al., 2023). These episodes, which triggered pronounced fluctuations in oil prices, are consistent with the historical patterns of oil shocks documented by Hamilton (2011).
Figure 1: Panel (a) displays the monthly log returns and panel (b) shows the squared monthly log returns of oil prices from January 1983 to December 2022 for six selected OPEC countries: Algeria, Iran, Libya, Nigeria, Saudi Arabia, and the United Arab Emirates.
Table LABEL:tab:stats reports the descriptive statistics for the monthly log returns of oil prices across the six major OPEC producers. All series exhibit positive means and medians, reflecting general upward trends but with varying dispersion levels, suggesting heterogeneous volatility. Negative skewness and high kurtosis are present in all cases, indicating asymmetric distributions with fat tails and substantial tail risk. These features—non-normality, asymmetry, and excess kurtosis—justify the use of heavy-tailed distributions in subsequent modeling.
Figure LABEL:fig:hist complements these findings by illustrating empirical distributions. While returns are broadly centered around zero, the dispersion and shape vary markedly. Libya stands out with the widest and most asymmetric distribution, followed by Nigeria, with a moderate spread. In contrast, Saudi Arabia and the UAE show tighter and more concentrated return profiles. All histograms deviate from normality, as confirmed by the Jarque and Bera (1980) tests in Table LABEL:tab:stats. These stylized facts are consistent with the well-documented empirical properties of financial returns, such as time-varying volatility and leptokurtosis (Bera and Higgins, 1993).
3.1 Interdependence of Oil Price Returns Among OPEC Countries
The correlation matrix in Table 1 shows strong positive comovement among the log returns of oil prices across the six OPEC countries, indicating high market integration and reflecting OPEC’s role in market stabilization. The highest correlations appear among North African countries, likely because of their shared economic structures and similar exposure to global shocks. Gulf countries also exhibit strong interdependence, consistent with their dominant position within OPEC. In contrast, slightly weaker correlations for Iran and the UAE may stem from geopolitical factors or divergent production policies. These findings suggest that while global shocks affect all members similarly, regional asymmetries persist, supporting the use of both constant and dynamic correlation models, such as CCC and DCC-GARCH (Bollerslev, 1990; Engle, 2002), as well as network-based approaches inspired by spatiotemporal ARCH models (see Otto et al., 2024b).
Table 1: Correlation Matrix of Monthly Oil Price Log Returns
Algeria Iran Libya Nigeria Saudi Arabia UAE
Algeria 1.0000 0.9518 0.9928 0.9946 0.9529 0.9376
Iran 0.9518 1.0000 0.9515 0.9525 0.9804 0.9677
Libya 0.9928 0.9515 1.0000 0.9934 0.9537 0.9350
Nigeria 0.9946 0.9525 0.9934 1.0000 0.9569 0.9356
Saudi Arabia 0.9529 0.9804 0.9537 0.9569 1.0000 0.9769
UAE 0.9376 0.9677 0.9350 0.9356 0.9769 1.0000
4 Overview of Time-Series Volatility Models
This section reviews the univariate ARCH and GARCH models, including their logarithmic variants (log-ARCH and log-GARCH), which provide the foundation for modeling time-varying volatility in financial time series. We then extend to multivariate specifications—CCC, DCC, and GO-GARCH—that capture static and dynamic correlations and volatility spillovers across assets, and are subsequently used to construct model-based weight matrices for the Network log-ARCH model.
4.1 Introduction to Univariate ARCH and GARCH-type Model
This subsection defines the univariate ARCH and GARCH models employed as building blocks in the CCC, DCC, and GO-GARCH frameworks. We also introduce the univariate log-ARCH process, which later serves as the foundation for the network-based log-volatility models.
4.1.1 Univariate ARCH and GARCH process
Since the seminal work of Engle (1982), autoregressive conditional heteroscedastic (ARCH) models have become central to econometric forecasting, allowing time-varying volatility to be explicitly modeled in financial time series. Bollerslev (1986) extended this framework through the Generalized ARCH (GARCH) model, where volatility depends on past squared observations and previous conditional variances. Both the ARCH and GARCH models capture conditional variance dynamics separately from the conditional mean, which may follow temporal trends, seasonal patterns, or autoregressive structures. In contrast, stochastic volatility models introduce a latent process to govern volatility evolution, which is also influenced by lagged squared returns. In what follows, we briefly review key univariate and multivariate versions of both classes. Let $\{Y_{t}\in\mathbb{R}^{r}:t\in\mathbb{Z}\}$ denote a discrete-time, $r$-dimensional random process observed at equally spaced intervals.
For a univariate time series (i.e., $r=1$), the process $Y_{t}$ is given by
$Y_{t}=h_{t}^{1/2}\varepsilon_{t}\,,$ (1)
where $h_{t}^{1/2}$ scales the i.i.d. sequence $\{\varepsilon_{t}\}$ with zero mean and unit variance, and $h_{t}$ denotes the conditional variance of $Y_{t}$, interpreted as time-varying volatility. In volatility models, $h_{t}$ evolves based on past realizations of the process.
In an ARCH($p$) model, the volatility dynamics are specified as
$h_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}^{2},$ (2)
while the GARCH($p,q$) specification is given by
$h_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}^{2}+\sum_{j=1}^{q}\beta_{j}h_{t-j},$ (3)
where $\alpha_{0}>0$, $\alpha_{i}\geq 0$, and $\beta_{j}\geq 0$ ensure the positivity of the conditional variance. The process is weakly stationary if $\sum_{i=1}^{p}\alpha_{i}+\sum_{j=1}^{q}\beta_{j}<1$. For a comprehensive treatment of GARCH-type models, see Francq and Zakoian (2019).
4.1.2 Log-ARCH and log-GARCH models
The log-volatility model was first proposed by Geweke (1986) and later supported by Pantula (1986) and Milhøj (1987). It avoids non-negativity constraints on the coefficients and facilitates the inclusion of regressors in the volatility specification. A log-ARCH model is given by
$\log h_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}\log Y_{t-i}^{2}.$ (4)
This can be extended to a log-GARCH model:
$\log h_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}\log Y_{t-i}^{2}+\sum_{j=1}^{q}\beta_{j}\log h_{t-j}.$ (5)
Log-GARCH models exhibit multiplicative volatility dynamics, unlike the additive structure of standard ARCH and GARCH models (see Francq et al., 2013). In these models, volatility is a deterministic function of past observations. In contrast, stochastic volatility (SV) models treat volatility as a latent process. A standard SV model specifies log-volatility as an AR(1) process:
$\log h_{t}-\mu_{h}=\phi(\log h_{t-1}-\mu_{h})+u_{t},$ (6)
where $|\phi|<1$, and $\varepsilon_{t}$ and $u_{t}$ are independent shocks. For details on the statistical properties of the SV models, see Ghysels et al. (1996) and Shephard (2005).
4.2 Multivariate GARCH models
While univariate GARCH models capture the conditional volatility of a single financial series, multivariate extensions are required to model the joint dynamics and cross-dependencies across assets. Bollerslev et al. (1988) introduced the VEC-GARCH model, which generalizes the GARCH framework to the multivariate setting by allowing time-varying covariance and correlation structures. Let $\mathbf{Y}_{t}$ be an $r$-dimensional random vector of asset returns. Then:
$\mathbf{Y}_{t}=\mathbf{H}_{t}^{1/2}\varepsilon_{t},$ (7)
where $\varepsilon_{t}$ is an i.i.d. innovation vector and $\mathbf{H}_{t}=\operatorname{Var}(\mathbf{Y}_{t}\mid\mathcal{F}_{t-1})$ is the conditional covariance matrix, with:
$\text{Var}(\mathbf{Y}_{t}\mid\mathbf{Y}_{v},v<t)=\mathbf{H}_{t}.$ (8)
Various multivariate GARCH models address the growing complexity of estimating $\mathbf{H}_{t}$ as the number of series increases. In this study, we focus on the CCC-GARCH, DCC-GARCH, and GO-GARCH models, each of which offers distinct features and levels of flexibility. These specifications are particularly suited for constructing model-based weight matrices in our network analysis of monthly log returns from the six OPEC countries. The following subsections discuss these models in increasing order of structural flexibility, highlighting their methodological linkages.
4.2.1 Constant and Dynamic Conditional Correlation GARCH Models
To capture co-movement in financial returns with time-varying volatility, Bollerslev (1990) proposed the Constant Conditional Correlation (CCC) GARCH model. This framework assumes that, while individual asset volatilities may evolve over time, their correlations remain constant. Formally, the conditional covariance matrix $\mathbf{H}_{t}$ is decomposed as:
$\mathbf{H}_{t}=\mathbf{D}_{t}\mathbf{R}\mathbf{D}_{t},$ (9)
where $\mathbf{R}$ is a fixed correlation matrix, and $\mathbf{D}_{t}$ is a diagonal matrix containing conditional standard deviations,
$\mathbf{D}_{t}=\text{diag}(\sqrt{h_{1t}},\ldots,\sqrt{h_{Nt}}).$ (10)
Each $h_{it}$ follows a univariate GARCH$(p,q)$ process, as defined in (3). Its general form is:
$h_{it}=\alpha_{0}^{(i)}+\sum_{j=1}^{p}\alpha_{j}^{(i)}Y_{i,t-j}^{2}+\sum_{j=1}^{q}\beta_{j}^{(i)}h_{i,t-j},\quad i=1,\ldots,N.$ (11)
While CCC-GARCH offers a simple yet effective structure, the assumption of constant correlations can be overly restrictive. To address this, Engle (2002) introduced the Dynamic Conditional Correlation (DCC) model, which preserves the same matrix form,
$\mathbf{H}_{t}=\mathbf{D}_{t}\mathbf{R}_{t}\mathbf{D}_{t},$ (12)
but allows the correlation matrix $\mathbf{R}_{t}$ to evolve over time. Correlations are derived from standardized residuals $\boldsymbol{\varepsilon}_{t}=\mathbf{D}_{t}^{-1}Y_{t}$, through a dynamic process:
$\mathbf{Q}_{t}=(1-a-b)\bar{\mathbf{Q}}+a\,\boldsymbol{\varepsilon}_{t-1}\boldsymbol{\varepsilon}_{t-1}^{\top}+b\,\mathbf{Q}_{t-1},$ (13)
where $\bar{\mathbf{Q}}$ is the unconditional covariance of $\boldsymbol{\varepsilon}_{t}$, and $a+b<1$ ensures the stationarity. The matrix $\mathbf{Q}_{t}$ is then rescaled to yield the time-varying correlation matrix:
$\mathbf{R}_{t}=\text{diag}(\mathbf{Q}_{t})^{-1/2}\mathbf{Q}_{t}\text{diag}(\mathbf{Q}_{t})^{-1/2}.$ (14)
By introducing dynamic correlations while preserving a parsimonious structure, the DCC model offers greater flexibility in modeling multivariate volatility dynamics. However, as the number of assets increases, the estimation of the correlation matrices becomes more complex. To circumvent this, the GO-GARCH model adopts a factor-based approach, as outlined below.
4.2.2 The Generalized Orthogonal GARCH Model
The Generalized Orthogonal GARCH (GO-GARCH) model, introduced by Van der Weide (2002), extends the Orthogonal GARCH framework by allowing the mixing matrix to be any invertible matrix, thus relaxing the orthogonality constraint. The $r$-dimensional zero-mean return vector ${Y}_{t}$ is expressed as a linear combination of uncorrelated latent factors ${f}_{t}$:
${Y}_{t}=\mathbf{Z}{f}_{t},\qquad{f}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{H}_{t}),$ (15)
where $\mathbf{Z}\in\mathbb{R}^{r\times r}$ is an invertible matrix and $\mathbf{H}_{t}=\text{diag}(h_{1t},\dots,h_{rt})$ is diagonal with each element following a univariate GARCH(1,1)11(1,1)( 1 , 1 ) process:
$h_{it}=\omega_{i}+\alpha_{i}f_{i,t-1}^{2}+\beta_{i}h_{i,t-1},\quad\text{for }i=1,\dots,r.$ (16)
The implied conditional covariance matrix of ${Y}_{t}$ is
$\mathbf{\Sigma}_{t}=\mathbf{Z}\mathbf{H}_{t}\mathbf{Z}^{\top}.$ (17)
The estimation proceeds in two steps. First, a spectral decomposition of the unconditional covariance matrix $\mathbf{H}=\mathbb{E}[{Y}_{t}{Y}_{t}^{\top}]$ provides initial estimates of the eigenvectors $\mathbf{P}$ and the eigenvalues $\mathbf{A}$. Second, a rotation matrix $\mathbf{U}$ is estimated via the maximum likelihood method, yielding:
$\mathbf{Z}=\mathbf{P}\mathbf{A}^{1/2}\mathbf{U}.$ (18)
The time-varying conditional correlation matrix is then recovered as
$\mathbf{R}_{t}=\mathbf{D}_{t}^{-1}\mathbf{\Sigma}_{t}\mathbf{D}_{t}^{-1},\qquad\mathbf{D}_{t}=\text{diag}(\sqrt{h_{1t}},\dots,\sqrt{h_{rt}}).$ (19)
The GO-GARCH model is particularly suitable for high-dimensional settings. By factorizing returns into latent components, it reduces the parameter space and avoids the direct estimation of full conditional covariance matrices. Each latent process follows a simple GARCH structure, making the model both flexible and parsimonious in capturing complex volatility dynamics, as in our network-based analysis of OPEC countries’ oil return volatilities.
5 Network-based Volatility models
A network is defined as $\mathcal{G}=(V,E)$, where $V$ denotes the set of vertices (e.g., OPEC countries) and $E$ the set of weighted edges capturing financial, geographic, or structural interdependencies; this structure underpins network-based log-ARCH models that embed both temporal and cross-sectional volatility dynamics to analyze and forecast spillovers.
5.1 Network log-ARCH models
The network log-ARCH model is an extension of spatiotemporal ARCH models (see Otto et al., 2018, 2024a), specifically designed to handle multivariate time series data in the context of network structures (Mattera and Otto, 2024). It captures both temporal and contemporaneous interactions in the volatility process, similar to the ARCH model, but adapted to a network setting. Unlike the previously reviewed univariate and multivariate GARCH models, these network-based approaches allow for volatility spillovers that occur instantaneously across the network in a GARCH-like manner. That is, the conditional variance in (8) also depends on adjacent nodes in $\mathbf{Y}_{t}$ (e.g., $v=t$), instead of strictly conditioning on all $v<t$. Moreover, multivariate GARCH models are typically not suitable for a large number of cross-sections. This is a significant extension, as it implies that the current volatility is directly influenced by the volatilities of the neighboring nodes.
In a network log-ARCH model, the process is observed at the nodes $V$, with the edges $E$ define the dependence structure in the volatility dynamics. This implies that the volatility at a given node is influenced not only by its own past observations but also by the volatility of its adjacent nodes at the same time point. Formally, the model builds upon the framework proposed by Otto et al. (2024a) for dynamic spatiotemporal ARCH models. The observed process $Y_{t}(s_{i})$ at time $t$ and node $v_{i}$ is given by
$Y_{t}(v_{i})=\sqrt{h_{t}(v_{i})}\varepsilon_{t}(v_{i}),$ (20)
where $h_{t}(v_{i})$ represents the volatility, and $\varepsilon_{t}(v_{i})$ is a white-noise process. In the network log-ARCH model, the volatility $h_{t}(v_{i})$ depends not only on past values of $Y_{t-1}(v_{i})$ but also on the contemporaneous values from adjacent nodes, $\{Y_{t}(v_{j}):j\in E_{i}\}$, where $E_{i}$ is the set of edges connected to node $v_{i}$.
Let $\boldsymbol{h}^{*}_{t}=(\ln h^{2}_{t}(v_{1}),\ldots,\ln h^{2}_{t}(v_{n}))^{\prime}$ and $\boldsymbol{Y}^{*}_{t}=(\ln Y^{2}_{t}(v_{1}),\ldots,\ln Y^{2}_{t}(v_{n}))^{\prime}$. The network log-ARCH process of order one can then be expressed as
$\boldsymbol{h}^{*}_{t}=\boldsymbol{\omega}+\mathbf{\Gamma}\boldsymbol{Y}^{*}_{t-1}+\rho\mathbf{W}\boldsymbol{Y}^{*}_{t},$ (21)
where $\mathbf{W}$ is the matrix of edge weights (adjacency matrix) defining the degree of volatility spillovers across the network, $\rho$ is a parameter capturing the strength of these contemporaneous network interactions, and $\mathbf{\Gamma}$ is a diagonal matrix of node-specific ARCH effects [^1] Note that this is a specific extension proposed by Mattera and Otto (2024), deviating from the original spatiotemporal log-ARCH model of Otto et al. (2024a).. The term $\boldsymbol{\omega}$ is constant volatility intercepts. The matrix $\mathbf{W}$, analogous to a spatial weight matrix, defines how the volatility spillovers propagate through the network.
The model can also be represented in an ARMA-like form of the log-squared observation:
$\boldsymbol{Y}^{*}_{t}=\boldsymbol{\phi}_{0}+\rho\mathbf{W}\boldsymbol{Y}^{*}_{t}+\mathbf{\Gamma}\boldsymbol{Y}^{*}_{t-1}+\boldsymbol{u}_{t},$ (22)
where $\boldsymbol{u}_{t}$ are the log-squared errors, and $\boldsymbol{\phi}_{0}$ is a constant vector. This formulation highlights the endogenous nature of the model, as $\boldsymbol{Y}^{*}_{t}=(\ln(Y^{2}_{t}(v_{i})))_{i\in V}$ appears on both sides of the equation, necessitating specialized estimation techniques such as the Generalized Method of Moments (GMM), as proposed by Otto et al. (2024a) or quasi-maximum-likelihood estimators (QMLE) by Otto et al. (2023) to effectively handle the endogeneity issue.
5.2 Choice of the Network Weight Matrices and Dissimilarity Measures
The choice of weight matrices $\mathbf{W}$ is based on the similarity between the log returns of the six OPEC countries and model-implied correlations, measured using temporal distance metrics. We consider six alternative dissimilarity measures: Euclidean distances, correlation-based distances, Piccolo distances, and distances derived from the CCC-GARCH, DCC-GARCH, and GO-GARCH models. Three of these were adopted from Mattera and Otto (2024), following the time series clustering methodologies outlined in Maharaj et al. (2019).
(1) Euclidean Distance. A raw-data-based approach computing pairwise distances between log-return observations:
$d_{ij}^{\text{(Euc)}}=\sqrt{\sum_{t=1}^{T}\left(y_{t}^{(i)}-y_{t}^{(j)}\right)^{2}},$ (23)
where $y_{t}^{(i)}$ denotes the oil price log return of country $i$ at time $t$. The corresponding network weight is defined as:
$W_{ij}^{\text{(Euc)}}=\frac{1}{d_{ij}^{\text{(Euc)}}},\quad W_{ii}^{\text{(Euc)}}=0,$ (24)
Therefore, smaller distances imply stronger connectivity.
(2) Correlation-Based Distance. This feature-based approach uses the Pearson correlation coefficient $\rho_{ij}^{\text{(corr)}}$ between countries $i$ and $j$, following the methodology of Mantegna (1999). The corresponding distance metric is:
$d_{ij}^{\text{(corr)}}=\sqrt{2(1-\rho_{ij}^{\text{(corr)}})},$ (25)
where $\rho_{ij}^{\text{(corr)}}$ is computed from the log-return series $\{y_{t}^{(i)}\}_{t=1}^{T}$ as:
$\rho_{ij}^{\text{(corr)}}=\frac{\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(i)}y_{t}^{(j)}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(i)}\right)\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(j)}\right)}{\sqrt{\left(\frac{1}{T}\sum_{t=1}^{T}(y_{t}^{(i)})^{2}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(i)}\right)^{2}\right)\left(\frac{1}{T}\sum_{t=1}^{T}(y_{t}^{(j)})^{2}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(j)}\right)^{2}\right)}}.$ (26)
This transformation maps high correlations ($\rho_{ij}^{\text{(corr)}}\approx 1$) to small distances ($d_{ij}^{\text{(corr)}}\approx 0$), providing a metric representation of similarity. The corresponding weight matrix is constructed as the inverse of the distance:
$W_{ij}^{\text{(Corr)}}=\frac{1}{d_{ij}^{\text{(corr)}}},\qquad W_{ii}^{\text{(Corr)}}=0,$ (27)
so that stronger co-movements translate into higher weights in the network structure.
(3) Piccolo-Based Distance. Following Piccolo (1990) and Mattera and Otto (2024), model-based dissimilarities in volatility dynamics are computed using the log-ARCH coefficient vectors from:
$\ln h_{t}^{(i)}=\omega_{i}+\sum_{p=1}^{P_{i}}\gamma_{ip}\ln(y_{t-p}^{(i)})^{2},$ (28)
where $\gamma_{ip}$ are the autoregressive coefficients for country $i$. Coefficient vectors are zero-padded for comparability when $P_{i}\neq P_{j}$. This corresponds to a truncated AR($\infty$) representation in the sense of Piccolo (1990). The Piccolo distance is defined as:
$d_{ij}^{\text{(Pic)}}=\sqrt{\sum_{p=1}^{P}\left(\gamma_{ip}-\gamma_{jp}\right)^{2}},$ (29)
with the associated weight matrix:
$W_{ij}^{\text{(Pic)}}=\frac{1}{d_{ij}^{\text{(Pic)}}},\quad W_{ii}^{\text{(Pic)}}=0.$ (30)
(4) CCC-GARCH Distance. Constant conditional correlations $\rho_{ij}^{\text{(CCC)}}$, extracted from the matrix $\mathbf{R}$ in (9) as introduced by Bollerslev (1990), are used to construct a dissimilarity measure:
$d_{ij}^{\text{(CCC)}}=\sqrt{2(1-\rho_{ij}^{\text{(CCC)}})}.$ (31)
To derive bounded and numerically stable network weights, we applied a nonlinear transformation of the squared distance:
$w_{ij}^{\text{(CCC)}}=\frac{1}{2(1-\rho_{ij}^{\text{(CCC)}})+1},\quad w_{ii}^{\text{(CCC)}}=0.$ (32)
This formulation ensures finite weights and prevents singularities when $\rho_{ij}^{\text{(CCC)}}\approx 1$, which corresponds to a near-perfect correlation.
(5) DCC-GARCH Distance. Based on the dynamic conditional correlations $\rho_{ij,t}^{\text{(DCC)}}$, which are elements of the time-varying matrix $\mathbf{R}_{t}$ in (14) (Engle, 2002), we compute time-averaged correlations:
$\bar{\rho}_{ij}^{\text{(DCC)}}=\frac{1}{T}\sum_{t=1}^{T}\rho_{ij,t}^{\text{(DCC)}},\quad d_{ij}^{\text{(DCC)}}=\sqrt{2\left(1-\bar{\rho}_{ij}^{\text{(DCC)}}\right)}.$ (33)
The corresponding network weights are defined as follows:
$w_{ij}^{\text{(DCC)}}=\frac{1}{2\left(1-\bar{\rho}_{ij}^{\text{(DCC)}}\right)+1},\quad w_{ii}^{\text{(DCC)}}=0.$ (34)
(6) GO-GARCH Distance. Based on the time-varying correlations $\rho_{ij,t}^{\text{(GO)}}$ extracted from the matrix $\mathbf{R}_{t}$ in (19), as proposed by Van der Weide (2002), we compute time-averaged correlations:
$\bar{\rho}_{ij}^{\text{(GO)}}=\frac{1}{T}\sum_{t=1}^{T}\rho_{ij,t}^{\text{(GO)}},\quad d_{ij}^{\text{(GO)}}=\sqrt{2(1-\bar{\rho}_{ij}^{\text{(GO)}})}.$ (35)
The associated network weights are as follows:
$W_{ij}^{\text{(GO)}}=\frac{1}{2(1-\bar{\rho}_{ij}^{\text{(GO)}})+1},\quad W_{ii}^{\text{(GO)}}=0.$ (36)
All raw spatial weight matrices—whether based on Euclidean, correlation, or GARCH-derived dissimilarities—are symmetric by construction, satisfying $W_{ij}=W_{ji}$ and $W_{ii}=0$, thereby ensuring well-defined pairwise comparisons. The GO-GARCH-based dissimilarities offer a complementary perspective to CCC and DCC specifications: while the latter directly model pairwise correlations, GO-GARCH captures time-varying dependence through latent factor structures. This factor-based approach enhances estimation stability in high-dimensional settings and yields smoother inter-country correlation dynamics for network construction.
6 Results and interpretation
This section presents the study’s empirical findings in three stages. First, we report and discuss the estimation results of the univariate and multivariate GARCH models, which form the basis for the construction of the network weight matrices. Second, we analyze the structural patterns revealed by the network plots, providing insight into the interdependencies among oil-exporting countries. Finally, we evaluate the out-of-sample forecasting performance of each modeling approach by comparing their root mean squared forecast errors (RMSFE) and mean absolute forecast errors (MAFE).
6.1 Model Estimation
As a foundation for the multivariate GARCH models employed in this study, we first estimate univariate GARCH(1,1) models assuming Student’s t-distributed innovations, with estimated degrees of freedom capturing moderate to strong tail heaviness. This distributional choice is motivated by the heavy-tailed nature of the monthly log-returns of oil prices, as evidenced by the initial analysis of the histograms presented in Figure LABEL:fig:hist.
Table 2: Univariate GARCH(1,1) estimates with Student’s t-innovations for monthly log-returns
Parameter Algeria Iran Libya Nigeria S. Arabia UAE
$\mu$ 0.0037 0.0069 0.0038 0.0029 0.0057 0.0055
(0.2781) (0.0333) (0.2842) (0.4446) (0.1383) (0.1114)
$\omega$ 0.0008 0.0014 0.0007 0.0007 0.0009 0.0008
(0.0766) (0.0270) (0.1689) (0.2630) (0.3305) (0.2962)
$\alpha_{1}$ 0.4432 0.5865 0.4496 0.4553 0.5391 0.5155
(0.0000) (0.0000) (0.0000) (0.0000) (0.0000) (0.0000)
$\beta_{1}$ 0.5408 0.4036 0.5494 0.5437 0.4599 0.4835
(0.0000) (0.0001) (0.0000) (0.0000) (0.0040) (0.0026)
$\nu$ 8.3936 5.4360 8.7382 8.8888 5.0903 5.1511
(0.0019) (0.0001) (0.0019) (0.0023) (0.0000) (0.0000)
LogLik 508.95 512.53 502.36 508.90 546.14 575.59
Notes: p-values are reported in parentheses. $\nu$ denotes the degrees of freedom of the Student’s t-distribution.
The univariate GARCH(1,1) estimates in Table 2 provide three key insights that motivate the transition to multivariate volatility modeling. First, all countries exhibit persistent volatility dynamics, with $\alpha_{1}+\beta_{1}\approx 1$, indicating long-lasting shock effects and stationary but highly persistent conditional variances. Second, the heterogeneity in the ARCH coefficients ($\alpha_{1}$) suggests differing short-run volatility responses, likely due to structural or informational asymmetries across markets. Third, the high log-likelihood values confirm the model’s adequacy in capturing volatility clustering and time-varying second moments. The estimated degrees of freedom for the Student’s t-distributed innovations further support the presence of heavy-tailed return distributions. Together, these findings justify the transition to multivariate GARCH models on three empirical grounds: (1) significant $\alpha_{1}$ and $\beta_{1}$ values confirm volatility clustering; (2) cross-country variation in parameters challenges the assumption of constant correlations; and (3) stationary yet persistent volatility ensures stable inputs for modeling interdependence. Hence, the univariate results offer both statistical and substantive foundations for estimating the CCC, DCC, and GO-GARCH models to examine cross-market volatility spillovers.
Table 3: Estimation Results: CCC-GARCH and DCC-GARCH with Multivariate t-Distribution
Parameter Algeria Iran Libya Nigeria Saudi Arabia UAE
Mean Equation
$\mu$ 0.0037 0.0069 0.0038 0.0029 0.0057 0.0055
(0.3717) (0.0533) (0.4565) (0.6401) (0.6069) (0.4984)
GARCH Equation
$\omega$ 0.0008 0.0014 0.0007 0.0007 0.0009 0.0008
(0.2293) (0.1278) (0.4298) (0.6195) (0.7925) (0.7263)
$\alpha$ 0.4432 0.5865 0.4496 0.4553 0.5391 0.5155
(0.00003) (0.00001) (0.00002) (0.0002) (0.0223) (0.0002)
$\beta$ 0.5408 0.4036 0.5494 0.5437 0.4599 0.4835
(0.0000) (0.0029) (0.0000) (0.0033) (0.3921) (0.2962)
Shape ($\nu$) 8.3941 5.4421 8.7422 8.8921 5.0938 5.1579
(0.0037) (0.00005) (0.0041) (0.0042) (0.0352) (0.0009)
DCC Parameters (for the DCC-GARCH)
$dcca_{1}$ 0.1016 (0.0000)
$dccb_{1}$ 0.8207 (0.0000)
Model Fit
Log-Likelihood 7486.96 (DCC)  /  7364.19 (CCC)
AIC -31.060 (DCC)  /  -30.556 (CCC)
BIC -30.642 (DCC)  /  -30.155 (CCC)
Notes: p-values are reported in parentheses below each coefficient estimate.
Table 3 reports the CCC-GARCH and DCC-GARCH estimation results for the monthly log returns of the six OPEC countries, assuming a multivariate Student’s t-distribution. The unconditional means are mostly positive but statistically insignificant, indicating no persistent abnormal returns. In the variance equations, both the ARCH ($\alpha$) and GARCH ($\beta$) coefficients are significant at the 1% level, confirming short- and long-term volatility persistence. The condition $\alpha+\beta<1$ holds across countries, ensuring covariance stationarity. The estimated degrees of freedom, ranging between 5 and 9, further support the appropriateness of modeling innovations using a heavy-tailed multivariate Student’s t-distribution. For correlation dynamics, the DCC parameters ($dcca_{1}=0.1016$, $dccb_{1}=0.8207$) are significant and satisfy $dcca_{1}+dccb_{1}<1$, indicating stable and persistent correlations. Compared to CCC-GARCH, the DCC-GARCH model yields higher log-likelihood and lower AIC/BIC values, supporting its superior fit in capturing the time-varying joint volatility structure across the OPEC markets.
Table 4: Estimation Results: GO-GARCH Model
Latent Factors (F1–F6)
Parameter F1 F2 F3 F4 F5 F6
Rotation Matrix $U$
Algeria -0.4324 0.3332 0.3640 -0.5228 -0.5299 -0.1250
Iran 0.2453 0.0099 0.6140 -0.4177 0.6103 0.1250
Libya 0.1985 -0.3691 0.6040 0.4155 -0.4553 0.2810
Nigeria -0.5830 0.2563 0.3190 0.5702 0.3451 -0.2210
Saudi Arabia 0.6094 0.6378 0.1110 0.2144 -0.1425 -0.3790
UAE 0.0461 -0.5294 0.1070 -0.0922 -0.0093 -0.8350
GARCH Parameters
$\omega$ 0.0681 0.1155 0.1545 0.1337 0.1434 0.1055
$\alpha_{1}$ 0.3776 0.2857 0.3056 0.8178 0.5916 0.2721
$\beta_{1}$ 0.6214 0.6175 0.5670 0.1812 0.3973 0.6490
Model Fit
Log-Likelihood 7351.5
AIC -30.62
BIC -30.46
Notes: The GO-GARCH model estimates latent factors (F1–F6) with constant mean and GARCH(1,1) volatility dynamics. The mixing matrix $A$ was omitted for brevity.
The GO-GARCH model offers a structural decomposition of volatility across OPEC countries by identifying the latent sources of co-movement. It supports a more parsimonious representation by modeling each latent factor with a univariate GARCH(1,1) process, thereby reducing dimensionality while preserving essential volatility dynamics. As shown in Table 4, the estimated rotation matrix $U$ reveals heterogeneous exposures of countries to different latent factors—some associated with short-term shocks and others with more persistent volatility—consistent with the covariance stationarity condition. This setup contrasts with the CCC and DCC models, which directly model the pairwise correlations. While DCC-GARCH flexibly captures time-varying dependence structures suitable for dynamic adjacency matrices, GO-GARCH provides stable and interpretable latent channels that inform the network’s underlying topology. Overall, GO-GARCH complements the DCC by enriching the network modeling framework with structural insights into systemic volatility transmission.
6.2 Network Construction
Following the estimation of the CCC-GARCH, DCC-GARCH, and GO-GARCH models, we constructed six distance-based matrices to support the network-based analysis. Two raw data–based matrices were computed using Euclidean and correlation-based dissimilarities from monthly log returns. Subsequently, model-based dissimilarity matrices were computed using truncated AR($\infty$) representations of log-squared returns, following Piccolo (1990), and form the conditional correlation structures produced by the CCC, DCC, and GO-GARCH models. All matrices were constructed following the methodology detailed in the previous section, ensuring comparability and coherence in the subsequent network-modeling framework.
Figure 2: Network plots of the six OPEC oil-exporting countries based on six different distance measures. Raw data-based distances: (a) Euclidean distance, (b) Correlation-based distance. Model-based distances: (c) Piccolo-based distance, (d) CCC-GARCH, (e) DCC-GARCH, and (f) GO-GARCH.
Figure 2 visualizes oil price interdependencies among the six OPEC countries through networks derived from monthly log returns. A notable contribution is the emergence of coherent and interpretable structures across all modeling approaches, which were constructed without relying on external variables. The raw-data-based networks—Euclidean (a) and correlation-based (b)—highlight regional proximity, with Algeria, Libya, and Nigeria forming a North African cluster, likely reflecting geographic and economic commonalities. In contrast, model-based networks reveal volatility-driven connections. The Piccolo-based distance (c), based on Piccolo (1990)’s AR($\infty$) structure, displays a different topology, suggesting the influence of long memory dynamics. The CCC-GARCH (d), DCC-GARCH (e), and GO-GARCH (f) networks exhibit a strong Middle Eastern core, with Saudi Arabia and the UAE emerging as central hubs, particularly Saudi Arabia, which appears as the primary source of risk spillovers to the majority of OPEC members, highlighting its clear capacity to propagate systemic risk among member countries. This dominant role aligns with its status as the largest oil producer within the organization (Fattouh, 2007; Alkhathlan et al., 2014; Al Rousan et al., 2018). Iran’s shifting position across models reflects external constraints, such as sanctions and production caps, which affect macroeconomic variables, including oil output, inflation, and government revenues (Nakhli et al., 2021). These networks, constructed solely from statistical features, highlight how spatiotemporal and GARCH-based models capture the latent interdependencies shaped by geopolitical and macroeconomic influences. These interdependencies may also reflect broader macroeconomic shocks and geopolitical frictions, as emphasized by Kilian (2008). Moreover, as noted by Mignon and Saadaoui (2024), such factors are essential for understanding systemic risk and volatility propagation within the global oil market. These visualizations identify dominant clusters and key players under varying model assumptions and serve as the structural foundation for the forecasting performance evaluation conducted in the next subsection.
6.3 Forecasting Results
First, we estimate the network-based volatility dynamics using the Generalized Method of Moments (GMM) framework proposed by Otto et al. (2024a) for dynamic spatiotemporal ARCH models. The models incorporate six different network weight matrices. Table 5 presents the estimation results for each network approach. Across all cases, we observe a pronounced instantaneous network ARCH effect ($\rho$), while the country-specific temporal ARCH effects ($\gamma_{1},\ldots,\gamma_{6}$) remain relatively small and partially statistically insignificant.
Table 5: GMM Estimation Results of the Dynamic Spatiotemporal ARCH Model under Different Network Structures
Network Weight Matrix $\rho$ $\gamma_{1}$ $\gamma_{2}$ $\gamma_{3}$ $\gamma_{4}$ $\gamma_{5}$ $\gamma_{6}$ $\sigma^{2}$
Euclidean-based 0.952 -0.016 0.027 -0.015 -0.130 0.202 0.219 1.498
Std. Error 0.054 0.045 0.042 0.049 0.056 0.042 0.047
t-statistic 17.545 -0.359 0.652 -0.315 -2.320 4.827 4.623
Correlation-based 0.955 -0.014 0.021 -0.022 -0.128 0.206 0.218 1.493
Std. Error 0.054 0.045 0.042 0.049 0.056 0.042 0.047
t-statistic 17.579 -0.301 0.503 -0.442 -2.294 4.956 4.615
Piccolo-based 1.213 -0.156 0.114 -0.149 -0.024 0.342 0.388 2.541
Std. Error 0.103 0.065 0.055 0.070 0.073 0.053 0.058
t-statistic 11.791 -2.402 2.094 -2.139 -0.323 6.414 6.645
CCC-GARCH-Based 0.923 0.003 0.004 0.007 -0.070 0.175 0.121 1.587
Std. Error 0.053 0.048 0.046 0.050 0.057 0.044 0.050
t-statistic 17.443 0.063 0.077 0.147 -1.231 3.980 2.393
DCC-GARCH-Based 0.923 0.003 0.004 0.007 -0.070 0.174 0.122 1.587
Std. Error 0.053 0.048 0.046 0.050 0.057 0.044 0.050
t-statistic 17.444 0.064 0.094 0.138 -1.238 3.968 2.408
GO-GARCH-Based 0.922 0.002 0.001 0.010 -0.067 0.175 0.124 1.598
Std. Error 0.053 0.048 0.046 0.050 0.057 0.044 0.051
t-statistic 17.378 0.039 0.016 0.193 -1.184 3.980 2.454
Notes: $\hat{\rho}$ is the spatial autoregressive coefficient; $\hat{\gamma}_{j}$ ($j=1,\ldots,6$) are the conditional variance parameters for each of the six OPEC countries. The six weight matrices are based on: Euclidean, Correlation, Piccolo, CCC-GARCH, DCC-GARCH, and GO-GARCH distances.
Following the estimation of the network-based volatility models, we assess their out-of-sample forecasting performance using two standard loss functions: root mean squared forecast error (RMSFE) and mean absolute forecast error (MAFE). These metrics are computed over the full forecast horizon $T$ and across all $N$ countries. Let $y_{i,t}$ denote the realized log-volatility and $\hat{y}_{i,t}$ its one-step-ahead forecast for country $i$ at time $t$. The average forecast errors are then given by:
$\mathrm{RMSFE}$ $=\frac{1}{N}\sum_{i=1}^{N}\sqrt{\frac{1}{T}\sum_{t=1}^{T}(y_{i,t}-\hat{y}_{i,t})^{2}},$ (37)
$\mathrm{MAFE}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{T}\sum_{t=1}^{T}|y_{i,t}-\hat{y}_{i,t}|.$ (38)
To formally test for differences in predictive accuracy, we employ the Diebold–Mariano (DM) test (Diebold and Mariano, 1995, 2002). Let $e^{(m)}_{i,t}=y_{i,t}-\hat{y}^{(m)}_{i,t}$ denote the forecast error from model $m$. The loss differential between models $m_{1}$ and $m_{2}$ is defined as:
$d_{i,t}=L(e^{(m_{1})}_{i,t})-L(e^{(m_{2})}_{i,t}),\qquad\text{with }L(e)=e^{2}\text{ or }|e|.$
The DM test statistic is computed as:
$\mathrm{DM}_{i}=\frac{\bar{d}_{i}}{\sqrt{\widehat{\mathrm{Var}}(\bar{d}_{i})}},\qquad\bar{d}_{i}=\frac{1}{T}\sum_{t=1}^{T}d_{i,t},$ (39)
where $\widehat{\mathrm{Var}}(\bar{d}_{i})$ denotes the heteroskedasticity and autocorrelation consistent (HAC) variance estimator of Newey and West (1987), with truncation lag $k-1=0$ for one-step-ahead forecasts. Two-sided $p$-values are reported at the 5% significance level.
For comparisons involving nested models, we employ the Clark–West (CW) test (Clark and West, 2007), which adjusts for the bias introduced by overfitting in the larger model’s mean squared prediction error (MSPE). The adjusted loss differential is defined as:
$f_{t}=(y_{t+\tau}-\hat{y}_{1,t+\tau})^{2}-\left[(y_{t+\tau}-\hat{y}_{2,t+\tau})^{2}-(\hat{y}_{1,t+\tau}-\hat{y}_{2,t+\tau})^{2}\right],$
where Model 2 nests Model 1. The CW test statistic is then computed as the average of the adjusted loss differentials:
$CW=\frac{1}{P}\sum_{t=R}^{T}f_{t},$ (40)
where $P=T-R+1$ is the number of out-of-sample observations, and $R$ denotes the initial estimation window. Under the null hypothesis of equal predictive accuracy, the test statistic is asymptotically normally distributed. One-sided critical values apply: the null is rejected at the 10% significance level if $CW>1.282$, and at the 5% level if $CW>1.645$.
Table 6: Forecast Inference Compared to Net-GO-GARCH: Bootstrap Confidence Intervals and Diebold and Mariano (2002) Tests (Sorted Worst to Best, Standard DCC-GARCH Excluded)
Model RMSFE CI MAFE CI DM $p$-value
Lower Upper Lower Upper
Std. GO-GARCH 5.7422 6.4045 5.3793 5.9524 $<0.001^{***}$
Piccolo-Diss 5.9414 6.6492 5.0637 5.6824 $0.0106^{**}$
Correlation-Diss 3.7825 4.7440 2.8536 3.5627 $<0.001^{***}$
Euclidean-Diss 3.5847 4.5018 2.7187 3.3796 $<0.001^{***}$
Std. CCC-GARCH 2.3063 2.9416 1.6726 2.1596 $0.0018^{***}$
Net. CCC-GARCH 2.1806 2.6795 1.6545 2.0391 $0.0811$
Net. DCC-GARCH 2.1806 2.6794 1.6547 2.0394 $0.1029$
Benchmark: Net-GO-GARCH 2.1797 2.6766 1.6541 2.0376
Notes: 95% bootstrap confidence intervals based on 2000 replications. Diebold–Mariano (DM) tests compare predictive accuracy against the benchmark network log-ARCH model with GO-GARCH-based distances. CW results were omitted to focus on non-nested comparisons. Significance levels: ${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$.
Table 7: Forecast Comparison: Best Network Models vs. Standard DCC-GARCH
Model RMSFE CI MAFE CI DM $p$ CW $p$
Lower Upper Lower Upper
BM: Std. DCC-GARCH 2.0844 2.7029 1.4689 1.9078
Net-GO 2.1797 2.6766 1.6541 2.0376 0.5793 0.6978
Net-DCC 2.1806 2.6794 1.6547 2.0394 0.5620 0.6981
Net-CCC 2.1806 2.6795 1.6545 2.0391 0.5621 0.6981
Std. CCC-GARCH 2.3063 2.9416 1.6726 2.1596 $<0.001^{***}$ 0.9771
Notes: Standard DCC-GARCH serves as the benchmark (BM). 95% bootstrap confidence intervals are based on 2000 replications. Predictive accuracy is evaluated using Diebold–Mariano (DM) and Clark–West (CW) tests. Network-based models do not differ significantly from the benchmark, whereas the standard CCC-GARCH model performs significantly worse. Significance levels: ${}^{***}p<0.001$.
The forecast evaluation highlights the comparative performance of the network-based log-ARCH models constructed using GARCH-informed distance-weight matrices. These models substantially reduce the computational burden relative to standard multivariate GARCH models, particularly during rolling window forecasting. As shown in Table 9, the network-based model incorporating GO-GARCH-based distances yields the lowest forecast errors relative to other network-based models as well as standard GO-GARCH and CCC-GARCH models under the baseline specification (RMSFE = 2.4182; MAFE = 1.8373), although standard DCC-GARCH remains the most accurate overall. These results were consistent across alternative training windows and zero-adjustment schemes. Its robustness suggests strong stability with respect to estimation perturbations. Comparable performance is observed for the log-ARCH models constructed using CCC-GARCH- and DCC-GARCH-based distances, with nearly identical baseline metrics and robust profiles. However, the standard DCC-GARCH remains the best among all specifications, with the clear advantage of being truly dynamic, while the network log-ARCH models with CCC, DCC, and GO-GARCH-based weight matrices use static time-averaged weight matrices. Nevertheless, they remain strong competitors. Their structure allows for a clear inclusion of instantaneous spillover effects, reduces computational burdens, and accounts for the interconnectedness of the OPEC members log returns.
The complementary inference results in Table 6 further support the statistical relevance of these models. While the static-distance-based variants (Euclidean, Correlation, Piccolo) perform significantly worse, the GARCH-informed networks (CCC, DCC, and GO) yield significantly more accurate forecasts than the standard GO-GARCH and CCC-GARCH benchmarks. The Diebold and Mariano (2002) test statistics indicate only marginal differences among the top three network configurations. In all pairwise comparisons between these models, the $p$-values associated with the DM test remained strictly above the 5% significance level, indicating that they are statistically indistinguishable from each other in terms of forecast performance.
Table 7 offers a focused comparison with the standard DCC-GARCH model. Although the standard DCC-GARCH in Table 9 achieves the lowest RMSFE (2.3854) and MAFE (1.6843) overall, its performance advantage over the best network-based models is modest and not statistically significant at conventional levels, in contrast to the standard CCC-GARCH model’s. This is corroborated by Table 7, which shows overlapping confidence intervals and $p$-values well above the 10% threshold across Diebold and Mariano (2002) and Clark and West (2007) test statistics for the GARCH-informed network-based log-ARCH models, confirming their statistical comparability to the standard DCC-GARCH specification.
Table 8: Model Confidence Set (MCS) Results ($\alpha=0.05$, $B=5000$, $T_{\text{max}}$)
Model RankM $v_{M}$ MCSM RankR $v_{R}$ MCSR Loss
Std. DCC-GARCH 1 -1.3295 1.0000 1 -0.3499 1.0000 5.7152
Net-GO-GARCH 2 -0.7799 1.0000 2 0.3499 0.9894 5.8637
Net-CCC-GARCH 3 -0.7477 1.0000 4 0.8501 0.8632 5.8709
Net-DCC-GARCH 4 -0.7475 1.0000 3 0.7895 0.8632 5.8710
Std. CCC-GARCH 5 1.4965 0.2088 5 2.1695 0.1056 6.8817
Note: Four models (Euclidean-Diss, Correlation-Diss, Piccolo-Diss, Standard GO-GARCH) were eliminated from the MCS. The standard DCC-GARCH achieved the lowest loss and remained superior under both $T_{\text{max}}$ rankings. Network models using CCC, DCC, and GO-GARCH weighting are statistically indistinguishable from the benchmark (5% level).
Table 8 reports the MCS outcomes at the 5% level based on the method of Hansen et al. (2011). The standard DCC-GARCH model achieves the lowest average forecast loss and ranks first under both the RMSFE and MAFE criteria. Importantly, the network log-ARCH models constructed from CCC, DCC, and GO-GARCH-based distance matrices are retained in the superior set, indicating a statistically equivalent forecasting performance. In contrast, models based on Euclidean, Correlation-based, and Piccolo dissimilarities were excluded because of significantly higher loss metrics. As illustrated in Figure 3, which displays the average forecast losses along with MCS inclusion indicators, these results reflect the advantage of incorporating volatility-driven dependencies. While DCC-GARCH benefits from a fully dynamic structure, even the standard GO- and CCC-GARCH models do not perform well in comparison. Despite the standard CCC-GARCH being retained in the MCS, it does not significantly outperform the other four alternatives at the 5% level than the four alternatives: the standard DCC-GARCH and the three network log-ARCH models with CCC, DCC, and GO-GARCH-based weights. This demonstrates the practical value of network-based models, which achieve forecasting accuracy comparable to fully dynamic specifications while using computationally efficient static weights.
Overall, the results demonstrate that network-based log-ARCH models constructed with GARCH-derived weight matrices offer a viable and interpretable alternative to standard multivariate GARCH specifications. In particular, the models based on GO-, DCC-, and CCC-GARCH distances, all deliver strong and robust out-of-sample performance. Despite using a time-invariant weight structure, these models remain competitive with the fully dynamic standard DCC-GARCH model. This finding suggests that the static, time-averaged GARCH-based weights effectively capture the persistent volatility relationships among OPEC members. Future research could explore whether incorporating time-varying network structures yields further improvements, particularly during periods of structural change in oil markets.
7 Conclusion and Outlook
This study introduced a novel application of GMM-based Network log-ARCH models to examine volatility spillovers and interconnections among oil-exporting countries within the OPEC framework. By leveraging six symmetric network weight matrices constructed from raw-data log returns, model-based dissimilarities, and time-averaged conditional correlation structures, we demonstrate how network-based volatility models can capture the complex transmission of risk within oil markets. Despite OPEC’s coordination efforts, the analysis reveals persistent heterogeneities across member countries, likely reflecting geopolitical frictions and structural asymmetries, as evidenced by the underlying network structures.
While the Network log-ARCH models incorporating CCC-, DCC-, and GO-GARCH-based distances exhibit robust forecasting performance, consistently outperform several standard multivariate GARCH alternatives, and reduce the computational burdens associated with these models, the standard DCC-GARCH model remains the most accurate overall. Its superiority is plausibly attributed to its ability to capture dynamic correlation structures in real time, whereas the network models rely on static, time-averaged weight matrices. This structural limitation points to a promising avenue for future research: the development of fully dynamic network-based log-ARCH models incorporating time-varying network weight matrices $W_{t}$. Such models could better reflect evolving transmission channels and further enhance forecasting performance in volatile environments. Additionally, ensemble forecasting methods, as proposed by Becker and Clements (2008), may complement these approaches and improve predictive robustness.
This study also highlights broader opportunities for extending the analysis. Expanding the framework to include all OPEC+ members—and eventually all major oil-producing and exporting countries—would yield a more comprehensive view of systemic risks in the global oil market. Moreover, incorporating explanatory variables related to geopolitical tensions, sanctions, and macroeconomic fundamentals (e.g., supply-demand imbalances) could enrich the model specification and improve its interpretability. Future extensions of this study may also explore machine learning techniques to enhance adaptability, capture nonlinear dynamics, and exploit historical regularities more effectively. These directions offer a compelling path toward improving volatility forecasting and advancing the understanding of risk transmission mechanisms in the global energy markets.
Appendix A Forecast Sensitivity Analysis
Table 9: Forecast-Error Magnitudes and Sensitivity Checks
Model Baseline Training Window ($T_{0}$) Zero-Adjustment ($\varepsilon$)
RMSFE MAFE 200 250 300 350 min 1%ile 1e-6
Network log-ARCH Models (with different weight matrices)
Euclidean-Diss 3.9511 3.0325 4.4192 4.1742 3.9511 4.1391 3.9511 3.8108 3.8230
Correlation-Diss 4.1527 3.1918 4.6287 4.3773 4.1527 4.3557 4.1527 4.0089 4.0215
Piccolo-Diss 5.9245 5.3727 6.0643 6.0301 5.9245 6.0394 5.9245 5.8645 5.8698
Net CCC-GARCH 2.4196 1.8374 2.8355 2.6929 2.4196 2.5000 2.4196 2.3355 2.3428
Net DCC-GARCH 2.4196 1.8374 2.8356 2.6929 2.4196 2.4999 2.4196 2.3355 2.3427
Net GO-GARCH 2.4182 1.8373 2.8338 2.6930 2.4182 2.4996 2.4182 2.3347 2.3419
Standard Multivariate GARCH Models
Std. CCC-GARCH 2.6184 1.9018 2.6858 2.7018 2.6184 2.7358 2.6184 2.5723 2.5760
Std. DCC-GARCH 2.3854 1.6843 2.5488 2.5013 2.3854 2.4896 2.3854 2.3363 2.3403
Std. GO-GARCH 6.0758 5.6669 6.2848 6.2449 6.0758 6.1759 6.0758 6.0462 6.0484
Notes: Baseline uses $T_{0}=300$, $\varepsilon=\min Y^{2}_{\neq 0}$. Bold entries highlight the best performance in both Network-based models and Standard Multivariate GARCH models.
Appendix B Model Losses and MCS Membership
Figure 3: Model Losses and MCS Membership
Note: Models eliminated from the MCS include the network-based log-ARCH models with Euclidean-, Correlation-, and Piccolo-based dissimilarities, as well as the standard GO-GARCH model. Green bars indicate models retained in the 95% Model Confidence Set (MCS); gray bars were eliminated. The standard DCC-GARCH model achieved the lowest forecast loss and remained in the superior set. Network models based on CCC, DCC, and GO-GARCH weights remain statistically indistinguishable from the benchmark at the 5% significance level.